OpenStudy (anonymous):

add the following polynomials \[(1/10x ^{2}-1/2x-3/10)+(1/2x ^{2}+1/4x-1/5)\]

OpenStudy (anonymous):

You can add the ones that are to the same power, such as 1/10x^2 and 1/2x^2 as well as -1/2x and 1/4x

OpenStudy (anonymous):

i dont get you

OpenStudy (anonymous):

1/10 + 1/2 =__ -1/2 + 1/4 =__ -3/10 + -1/5 = __

OpenStudy (anonymous):

so 0.6 -0.25 -1.6

OpenStudy (anonymous):

Now just plug in the variables next to each one.

OpenStudy (anonymous):

did i do it correct because am not sure what it is you are asking

OpenStudy (anonymous):

.6x^2-.25x-1.6 would be your final answer if you added everything correctly.